# Dynamic Scene Understanding through Object-Centric Voxelization and Neural Rendering

### v1

@@ 4.3 Scene Decomposition
[…]
@@
$\delta_{\text{den}}=1,\alpha_{p}=0.01$ (Final model) 32.58 90.60 31.78 85.41 28.82 92.61
4.4 Dynamic Scene Editing
After the training period, the object-centric voxel representations learned by DynaVol-S can be readily used in downstream tasks such as scene editing without the need for additional model tuning. DynaVol-S allows for easy manipulation of the observed scene by directly modifying the object occupancy values within the voxel grids or switching the learned deformation function to a pre-defined one. This flexibility empowers users to make various scene edits and modifications. For instance, in the first example in Fig. 7, we remove the hand that is pinching the toys. In the second example, we modify the dynamics of the shoe from falling to rotating. More showcases are included in the Appendix.
4.5 Further Analyses
Ablation studies. We present ablation study results in Table VI. First, we can find that the forward deformation network $f_{\xi}^{\prime}(\cdot)$ is crucial to the final rendering and decomposition results. Furthermore, in the absence of the multi-grids joint optimization stage, the performance of DynaVol-S significantly degrades, highlighting the importance of refining the object-centric voxel representation with a latent-code-based renderer. Additionally, we conduct an ablation study that performs multi-grids joint optimization from scratch with randomly initialized $\mathcal{V}^{\text{M}}$ and $f_{\phi}(\cdot)$. The results clearly show that excluding the warmup stage substantially impacts the final performance, especially for the scene decomposition results.
The number of object-centric latent codes. First, we examine the impact of the number of latent codes ($N$), which indicates the maximum number of objects in the scene. According to Table VII, the model achieves comparable results as long as $N$ exceeds the number of objects in the scene. The presence of redundant latent codes has only a minor influence on both rendering and decomposition results. Additionally, we showcase the per-object rendering results in Appendix Fig. 14, demonstrating DynaVol-S’s robustness to variations in the number of latent codes. Our findings reveal that redundant latent codes typically remain unassigned, effectively addressing the potential issue of over-segmentation.
@@ 4.5 Further Analyses
Hyperparameters for the loss functions. The hyperparameters in the loss functions are set to $\alpha_{p}=0.01$, $\alpha_{b}=0.01$, $\alpha_{e}=0.008$, $\alpha_{c}=1.0$. We study the hyperparameter choice of the threshold value for splitting the foreground and background in the warmup stage, and the weight ($\alpha_{p}$) of per-point RGB loss in Table VIII. It can be found that our model is robust to the threshold, and the performance of DynaVol-S remains largely unaffected with different thresholds. Furthermore, it can be found that $\alpha_{p}=0.01$ performs better than $\alpha_{p}=0.0$, indicating that a conservative use of $\mathcal{L}_{\text{point}}$ can improve the performance. A possible reason is that it eases the training process by moderately penalizing the discrepancy of nearby sampling points on the same ray. When the weight of $\mathcal{L}_{\text{point}}$ increases to $\alpha_{p}=0.1$, the performance of the method decreases significantly. Such a substantial emphasis on $\mathcal{L}_{\text{point}}$ is intuitively unreasonable and can potentially introduce bias to the neural rendering process, thereby negatively impacting the final results. For the entropy loss in Eq. (10), it encourages the semantic probabilities $\widehat{\mathcal{A}}$ learned in the warmup phase to approach a one-hot distribution. In Fig. 8, we visualize $\widehat{\mathcal{A}}$ learned with different values of $\alpha_{e}$. By increasing the weight of entropy loss $\alpha_{e}$, we prevent the model from over-segmenting the 3D scene with a redundant number of slot features. In other words, tuning $\alpha_{e}$ allows us to use a larger default number of $N$ while balancing between over-segmentation and under-segmentation.
@@
TABLE IX: The analysis of using different semantic feature extractors.
@@
Alternative designs for volume slot attention. In considering the design of the volume slot attention module, we perform the experiments shown in Fig. 9. A straightforward approach, referred to as “Baseline”, rewrites Eq. (5) by removing the slot attention operator, such that: $\mathbf{f}_{i}^{s}=\text{Interp}\left(\mathbf{x}_{i}+f_{\psi}(\mathbf{x}_{i},t),\mathcal{V}^{\text{S}}\right)$ and $\widehat{F}(\mathbf{r})=\sum_{i=1}^{P}T_{i}\left(1-\exp(-\sigma_{i}\delta_{i})\right)\mathbf{f}_{i}^{s}$. In the Baseline model, we otherwise obtain $\mathcal{A}_{i}$ by applying the connected components algorithm to the cosine similarities greater than $0.9$ between adjacent voxel features. We can find that the Baseline model tends to merge features into a single cluster in $\widehat{\mathcal{A}}$. As shown, the ambiguous estimation of $\widehat{\mathcal{A}}$ results in a noisy outcome for the learned semantic features, thereby undermining the final scene decomposition results presented by $\mathcal{V}^{M}$. Another alternative approach uses the Gumbel-softmax to replace the original softmax operation in Eq. (5), which ensures that each voxel is exclusively associated with a single slot, with $\mathcal{A}_{i}$ representing a one-hot vector. In Fig. 9, we visualize the generated semantic features $\widehat{F}(\mathbf{r})$, the corresponding semantic probabilities $\widehat{\mathcal{A}}(\mathbf{r})$ obtained from Eq. (6), and the scene decomposition results $\widehat{P}(\mathbf{r})$. As we can see, the Gumbel-softmax method leads to noisy decomposition results due to its hard assignment strategy. Our method, in contrast, effectively mitigates the overfitting of 2D semantic feature noise, resulting in clearer boundaries and more accurate decomposition outcomes. It is important to note that $\mathcal{A}_{i}$ plays a crucial role in initializing 4D voxel grids, $\mathcal{V}^{\text{M}}$, which in turn affects the final decomposition results. Specifically, if $\mathcal{A}_{i}$ merges features excessively, DynaVol-S will lose the insights provided by the semantic features, potentially degrading to DynaVol.
@@
Alternative semantic feature extractors. In addition to DINOv2, our approach can be seamlessly integrated with other semantic feature extractors, such as MaskClip [34], ViT [35], and DINOv1 [32]. In Table IX, we explore the impact of these models. The results show that DINOv2 outperforms the other models. Aside from feature quality, these discrepancies may also stem from DINOv2’s strong capabilities in maintaining multi-view consistency [36].

### v2

@@ 4.5 Model Analyses
PSNR mIOU PSNR mIOU PSNR mIOU
@@
Final model 32.58 90.60 31.78 85.41 28.82 92.61
@@
Hyperparameters for the loss functions. The hyperparameters in the loss functions are set to $\alpha_{p}=0.01$, $\alpha_{b}=0.01$, $\alpha_{e}=0.008$, and $\alpha_{c}=1.0$. We study the hyperparameter choice of the threshold value for splitting the foreground and background in the warmup stage, and the weight ($\alpha_{p}$) of per-point RGB loss in Table IX. We find that the performance of DynaVol-S remains robust to different threshold values. Furthermore, $\alpha_{p}=0.01$ performs better than $\alpha_{p}=0.0$, indicating that a conservative use of $\mathcal{L}_{\text{point}}$ can improve the performance. A possible reason is that it eases the training process by moderately penalizing the discrepancy of nearby sampling points on the same ray. When the weight of $\mathcal{L}_{\text{point}}$ increases to $\alpha_{p}=0.1$, the method’s performance decreases significantly. It is because too much emphasis on $\mathcal{L}_{\text{point}}$ is intuitively unreasonable and can potentially introduce bias to the neural rendering process. The entropy loss in Eq. (10) encourages the semantic probabilities $\widehat{\mathcal{A}}$ learned in the warmup phase to approach a one-hot distribution. In Fig. 8, we visualize $\widehat{\mathcal{A}}$ learned with different values of $\alpha_{e}$. By increasing $\alpha_{e}$, we prevent the model from over-segmenting the 3D scene with a redundant number of slot features. In other words, tuning $\alpha_{e}$ allows us to use a larger default number of $N$ while balancing between over-segmentation and under-segmentation.
Figure 10: Visualization of per-object mesh reconstruction results on 3ObjRealSimp, with each object denoted by a different color.
TABLE X: The analysis of using different semantic feature extractors.
@@
TABLE XI: The analysis of model efficiency on real-world datasets and resource requirements of each training stage of our model.
Method Train (hour) Memory (GB) Test (FPS)
HyperNeRF 32 128 0.04
$\text{D}^{2}$eRF 8 320 0.04
TineuVox 0.5 22 0.22
DynaVol 3 16 0.25
DynaVol-S 1.8 21 0.28
Stage 1 1 21 N/A
Stage 2 0.3 N/A N/A
Stage 3 0.5 14 N/A
TABLE XII: Analyses of different training strategies for the bi-directional deformation networks.
Method 3ObjRand 8ObjFall 3ObjRealCmpx Chicken Average
[…]
Separate Training 32.45 90.60 31.92 86.43 28.86 92.61 28.44 90.55 30.41 90.05
Co-Training 32.58 90.60 31.78 85.41 28.82 92.61 28.43 90.70 30.40 89.83
Stop-gradient 32.23 80.34 30.32 69.63 28.77 92.60 28.44 90.74 29.94 83.33
@@
Alternative designs for volume slot attention. In considering the design of the volume slot attention module, we perform the experiments shown in Fig. 9. A straightforward approach, referred to as “Baseline”, rewrites Eq. (5) by removing the slot attention operator, such that: $\mathbf{f}_{i}^{s}=\text{Interp}\left(\mathbf{x}_{i}+f_{\psi}(\mathbf{x}_{i},t),{\mathcal{V}^{\text{Sem}}}\right)$ and $\widehat{F}(\mathbf{r})=\sum_{i=1}^{P}T_{i}\left(1-\exp(-\sigma_{i}\delta_{i})\right)\mathbf{f}_{i}^{s}$. In the Baseline model, we otherwise obtain $\mathcal{A}_{i}$ by applying the connected components algorithm to the cosine similarities greater than $0.9$ between adjacent voxel features. We can find that the Baseline model tends to merge features into a single cluster in $\widehat{\mathcal{A}}$. As shown, the ambiguous estimation of $\widehat{\mathcal{A}}$ results in a noisy outcome for the learned semantic features, thereby undermining the final scene decomposition results presented by $\mathcal{V}^{\text{Occ}}$. Another alternative approach uses the Gumbel-softmax to replace the original softmax operation in Eq. (5), which ensures that each voxel is exclusively associated with a single slot, with $\mathcal{A}_{i}$ representing a one-hot vector. In Fig. 9, we visualize the generated semantic features $\widehat{F}(\mathbf{r})$, the corresponding semantic probabilities $\widehat{\mathcal{A}}(\mathbf{r})$ obtained from Eq. (6), and the scene decomposition results $\widehat{P}(\mathbf{r})$. As we can see, the Gumbel-softmax method leads to noisy decomposition results due to its hard assignment strategy. Our method, in contrast, effectively mitigates the overfitting of 2D semantic feature noise, resulting in clearer boundaries and more accurate decomposition outcomes. It is important to note that $\mathcal{A}_{i}$ plays a crucial role in initializing 4D voxel grids, ${\mathcal{V}^{\text{Occ}}}$, which in turn affects the final decomposition results. Specifically, if $\mathcal{A}_{i}$ merges features excessively, DynaVol-S will lose the insights provided by the semantic features, potentially degrading to DynaVol.
@@
Alternative semantic feature extractors. In addition to DINOv2, our approach can be seamlessly integrated with other semantic feature extractors, such as MaskClip [34], ViT [35], and DINOv1 [32]. In Table X, we explore the impact of these models. The results show that DINOv2 outperforms the other models. Aside from feature quality, these discrepancies may also stem from DINOv2’s strong capabilities in maintaining multi-view consistency [36]. In the supplementary materials, we further explore the impact of using different variants of the DINOv2 model on the final performance of 3D decomposition, particularly for objects with intricate shapes.
Visualized per-object mesh reconstruction. To provide an intuitive understanding of its decomposition performance, we use the learned occupancy field to reconstruct the mesh for each object. As shown in Fig. 10, DynaVol-S demonstrates significant improvements over the baselines, including MovingParts, in both the accuracy of reconstructed geometries and decomposition results.
Model efficiency. In Table XI, we compare DynaVol-S with existing approaches in terms of training time, memory consumption, and rendering FPS on real-world data. Additionally, we provide a breakdown of the resource requirements for each training stage of our model. The results demonstrate that DynaVol-S achieves comparable training efficiency to the baseline models. Furthermore, due to our explicit opacity representation, DynaVol-S significantly reduces computational costs by filtering out low-opacity points, leading to the highest rendering FPS and reduced memory usage across all evaluated methods. Although DynaVol-S requires a longer training time than TineuVox, the primary overhead arises from training the occupancy and semantic voxel grids, $\mathcal{V}^{\text{Occ}}$ and $\mathcal{V}^{\text{Sem}}$, which facilitate scene decomposition—a capability absent in other methods. A potential direction for future research is to explore the use of K-Planes [37] to further minimize training time and memory consumption.
Analyses of bi-directional deformation networks. As discussed in Sec. 3.4.1, training the forward and the backward deformation networks simultaneously may raise concerns about whether both networks can be properly trained. To address this issue, we implement two alternative training strategies: “Separate” and “Stop-gradient”. In the separate strategy, we first use the image reconstruction loss to train the backward deformation network $f_{\psi}(\cdot)$, then use the fixed $f_{\psi}(\cdot)$ to supervise the forward modeling network $f^{\prime}_{\xi}(\cdot)$, ensuring both networks are well-initialized before joint training. The “Stop-gradient” strategy involves applying stop gradients to either side of Eq. (9), such that:
$\begin{split}\mathcal{L}_{\text{Cyc}}=\frac{1}{|\mathcal{R}|}\sum_{{r}\in\mathcal{R}}\frac{1}{P}\sum_{i=1}^{P}\Big{(}&\left\|\mathrm{sg}(f_{\psi}(x_{i},t))+f^{\prime}_{\xi}(x^{\prime}_{i},t)\right\|_{2}^{2}\\ +&\left\|f_{\psi}(x_{i},t)+\mathrm{sg}(f^{\prime}_{\xi}(x^{\prime}_{i},t))\right\|_{2}^{2}\Big{)}.\end{split}$ (12)
As shown in Table XII, the separate training method yields comparable performance to the proposed co-training method, while the stop-gradient method degrades model performance, particularly in synthetic scenes. We adopt the co-training strategy in DynaVol-S for its balance of performance and simplicity.